# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ Appendix E Reinforcement Learning Generated Subgraph Details
L3 $0233\scriptstyle\pm 0.0252 0.0$$020\scriptstyle\pm 0.0020 0.0$$035\scriptstyle\pm 0.0033 0.0$$018\scriptstyle\pm 0.0017 0.0$$133\scriptstyle\pm 0.0231 0.0$$0067\scriptstyle\pm 0.0115$
L3 + KG $0055\scriptstyle\pm 0.0055 0.0$$033\scriptstyle\pm 0.0031 0.0$$041\scriptstyle\pm 0.0039 0.0$$21\scriptstyle\pm 0.0020 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) $0300\scriptstyle\pm 0.0249 0.0$$220\scriptstyle\pm 0.0278 0.0$$199\scriptstyle\pm 0.0214 0.0$$104\scriptstyle\pm 0.0114 0.0$$267\scriptstyle\pm 0.0462 0.0$$0333\scriptstyle\pm 0.0306$
L3-FT(Med) + KG $0044\scriptstyle\pm 0.0038 0.0$$033\scriptstyle\pm 0.0031 0.0$$032\scriptstyle\pm 0.0028 0.0$$16\scriptstyle\pm 0.0014 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5011\scriptstyle\pm 0.0335$ $0.4727\scriptstyle\pm 0.0514$ $0.4862\scriptstyle\pm 0.0430$ $0.3230\scriptstyle\pm 0.0384$ $0.8400\scriptstyle\pm 0.0800$ $0.8267\scriptstyle\pm 0.0808$
L3-FT(QA) + KG $0.5116\scriptstyle\pm 0.0216$ $0.4807\scriptstyle\pm 0.0469$ $0.4950\scriptstyle\pm 0.0351$ $0.3301\scriptstyle\pm 0.0312$ $0.8800\scriptstyle\pm 0.0000$ $0.8333\scriptstyle\pm 0.0231$
G-Retriever $0.4417\scriptstyle\pm 0.0005$ $0.3680\scriptstyle\pm 0.0104$ $0.3995\scriptstyle\pm 0.0073$ $0.2504\scriptstyle\pm 0.0062$ $0.7200\scriptstyle\pm 0.0000$ $0.7667\scriptstyle\pm 0.0115$
GALAX $\mathbf{0.5272}\scriptstyle\pm 0.0143$ $\mathbf{0.5120}\scriptstyle\pm 0.0087$ $\mathbf{0.5192}\scriptstyle\pm 0.0103$ $\mathbf{0.3512}\scriptstyle\pm 0.0098$ $\mathbf{0.9200}\scriptstyle\pm 0.0693$ $\mathbf{0.8400}\scriptstyle\pm 0.0200$
Table 12: Model performances on PAAD
@@
L3 $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + KG $139\scriptstyle\pm 0.0241 0.000$$08\scriptstyle\pm 0.0014 0.0$$16\scriptstyle\pm 0.0027 0.000$$08\scriptstyle\pm 0.0014 0.0$$167\scriptstyle\pm 0.0289 0.0$$0083\scriptstyle\pm 0.0144$
L3-FT(Med) $0051\scriptstyle\pm 0.0061 0.0$$050\scriptstyle\pm 0.0043 0.0$$039\scriptstyle\pm 0.0036 0.0$$020\scriptstyle\pm 0.0018 0.0$$167\scriptstyle\pm 0.0289 0.0$$0083\scriptstyle\pm 0.0144$
L3-FT(Med) + KG $005\scriptstyle\pm 0.0008 0.000$$8\scriptstyle\pm 0.0014 0.000$$6\scriptstyle\pm 0.0010 0.000$$3\scriptstyle\pm 0.0005 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(QA) $0.5649\scriptstyle\pm 0.0559$ $0.5358\scriptstyle\pm 0.0686$ $0.5492\scriptstyle\pm 0.0605$ $0.3811\scriptstyle\pm 0.0584$ $0.8000\scriptstyle\pm 0.0500$ $0.8417\scriptstyle\pm 0.0577$
L3-FT(QA) + KG $0.5499\scriptstyle\pm 0.0570$ $0.5183\scriptstyle\pm 0.0794$ $0.5330\scriptstyle\pm 0.0683$ $0.3657\scriptstyle\pm 0.0650$ $0.8333\scriptstyle\pm 0.0764$ $0.8333\scriptstyle\pm 0.0520$
G-Retriever $0.5594\scriptstyle\pm 0.0114$ $0.4567\scriptstyle\pm 0.0058$ $0.5016\scriptstyle\pm 0.0009$ $0.3387\scriptstyle\pm 0.0007$ $\mathbf{0.8500}\scriptstyle\pm 0.0000$ $\mathbf{0.8500}\scriptstyle\pm 0.0000$
GALAX $\mathbf{0.5771}\scriptstyle\pm 0.0174$ $\mathbf{0.5708}\scriptstyle\pm 0.0118$ $\mathbf{0.5739}\scriptstyle\pm 0.0145$ $\mathbf{0.4044}\scriptstyle\pm 0.0154$ $0.8333\scriptstyle\pm 0.0289$ $\mathbf{0.8500}\scriptstyle\pm 0.0000$
Table 13: Model performances on GBM
@@
L3 $0056\scriptstyle\pm 0.0096 0.0$$011\scriptstyle\pm 0.0019 0.0$$19\scriptstyle\pm 0.0032 0.000$$9\scriptstyle\pm 0.0016 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) $0069\scriptstyle\pm 0.0045 0.0$$133\scriptstyle\pm 0.0088 0.0$$090\scriptstyle\pm 0.0059 0.0$$46\scriptstyle\pm 0.0030 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0107\scriptstyle\pm 0.0158 0.0$$056\scriptstyle\pm 0.0038 0.0$$064\scriptstyle\pm 0.0074 0.0$$32\scriptstyle\pm 0.0037 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(QA) $\mathbf{0.5641}\scriptstyle\pm 0.0310$ $0.5333\scriptstyle\pm 0.0462$ $\mathbf{0.5480}\scriptstyle\pm 0.0387$ $\mathbf{0.3794}\scriptstyle\pm 0.0364$ $0.9333\scriptstyle\pm 0.0667$ $\mathbf{0.9000}\scriptstyle\pm 0.0333$
L3-FT(QA) + KG $0.5391\scriptstyle\pm 0.0213$ $0.5178\scriptstyle\pm 0.0164$ $0.5281\scriptstyle\pm 0.0181$ $0.3596\scriptstyle\pm 0.0165$ $\mathbf{0.9556}\scriptstyle\pm 0.0385$ $0.8444\scriptstyle\pm 0.0694$
G-Retriever $0.4906\scriptstyle\pm 0.0075$ $0.4411\scriptstyle\pm 0.0038$ $0.4642\scriptstyle\pm 0.0010$ $0.3037\scriptstyle\pm 0.0000$ $\mathbf{0.9556}\scriptstyle\pm 0.0385$ $0.8333\scriptstyle\pm 0.0577$
GALAX $0.5527\scriptstyle\pm 0.0021$ $\mathbf{0.5422}\scriptstyle\pm 0.0139$ $0.5473\scriptstyle\pm 0.0082$ $0.3769\scriptstyle\pm 0.0079$ $0.9111\scriptstyle\pm 0.0770$ $0.8667\scriptstyle\pm 0.0000$
Table 14: Model performances on SARC
@@
L3 $0111\scriptstyle\pm 0.0192 0.0$$033\scriptstyle\pm 0.0058 0.0$$051\scriptstyle\pm 0.0089 0.0$$26\scriptstyle\pm 0.0045 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
@@
L3-FT(Med) $0689\scriptstyle\pm 0.0631 0.0$$167\scriptstyle\pm 0.0115 0.0$$229\scriptstyle\pm 0.0180 0.0$$119\scriptstyle\pm 0.0094 0.0$$222\scriptstyle\pm 0.0385 0.0$$0667\scriptstyle\pm 0.0577$
@@
L3-FT(QA) $0.5347\scriptstyle\pm 0.0167$ $0.5089\scriptstyle\pm 0.0353$ $0.5211\scriptstyle\pm 0.0266$ $0.3547\scriptstyle\pm 0.0253$ $\mathbf{0.9111}\scriptstyle\pm\mathbf{0.0385}$ $0.8778\scriptstyle\pm 0.0192$
L3-FT(QA) + KG $\mathbf{0.5423}\scriptstyle\pm\mathbf{0.0207}$ $0.5022\scriptstyle\pm 0.0539$ $0.5207\scriptstyle\pm 0.0383$ $0.3570\scriptstyle\pm 0.0344$ $0.8444\scriptstyle\pm 0.0385$ $0.8444\scriptstyle\pm 0.0385$
G-Retriever $0.4461\scriptstyle\pm 0.0308$ $0.3667\scriptstyle\pm 0.0115$ $0.4009\scriptstyle\pm 0.0055$ $0.2518\scriptstyle\pm 0.0051$ $0.8667\scriptstyle\pm 0.0000$ $\mathbf{0.9222}\scriptstyle\pm\mathbf{0.0192}$
GALAX $0.5414\scriptstyle\pm 0.0143$ $\mathbf{0.5144}\scriptstyle\pm\mathbf{0.0168}$ $\mathbf{0.5271}\scriptstyle\pm\mathbf{0.0157}$ $\mathbf{0.3602}\scriptstyle\pm\mathbf{0.0165}$ $\mathbf{0.9111}\scriptstyle\pm\mathbf{0.1018}$ $0.8444\scriptstyle\pm 0.0509$
Table 15: Model performances on OV
@@
L3 $0241\scriptstyle\pm 0.0251 0.0$$044\scriptstyle\pm 0.0051 0.0$$075\scriptstyle\pm 0.0085 0.0$$38\scriptstyle\pm 0.0043 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3 + KG $0660\scriptstyle\pm 0.0667 0.0$$111\scriptstyle\pm 0.0102 0.0$$183\scriptstyle\pm 0.0161 0.0$$93\scriptstyle\pm 0.0082 0.000$$00\scriptstyle\pm 0.0000 0.0$$0889\scriptstyle\pm 0.0770$
L3-FT(Med) $0356\scriptstyle\pm 0.0412 0.0$$100\scriptstyle\pm 0.0067 0.0$$102\scriptstyle\pm 0.0072 0.0$$051\scriptstyle\pm 0.0036 0.0$$444\scriptstyle\pm 0.0770 0.0$$0222\scriptstyle\pm 0.0385$
L3-FT(Med) + KG $0797\scriptstyle\pm 0.0496 0.0$$344\scriptstyle\pm 0.0280 0.0$$457\scriptstyle\pm 0.0375 0.0$$253\scriptstyle\pm 0.0226 0.0$$000\scriptstyle\pm 0.0000 0.0$$0333\scriptstyle\pm 0.0333$
L3-FT(QA) $0.5469\scriptstyle\pm 0.0274$ $0.5167\scriptstyle\pm 0.0406$ $0.5307\scriptstyle\pm 0.0325$ $0.3631\scriptstyle\pm 0.0295$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9111\scriptstyle\pm 0.0694$
L3-FT(QA) + KG $0.5439\scriptstyle\pm 0.0066$ $0.4989\scriptstyle\pm 0.0195$ $0.5200\scriptstyle\pm 0.0123$ $0.3542\scriptstyle\pm 0.0095$ $0.9556\scriptstyle\pm 0.0385$ $0.9444\scriptstyle\pm 0.0385$
G-Retriever $0.5413\scriptstyle\pm 0.0525$ $0.4033\scriptstyle\pm 0.0173$ $0.4575\scriptstyle\pm 0.0067$ $0.2991\scriptstyle\pm 0.0041$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0000}$
GALAX $\mathbf{0.5489}\scriptstyle\pm\mathbf{0.0176}$ $\mathbf{0.5344}\scriptstyle\pm\mathbf{0.0259}$ $\mathbf{0.5413}\scriptstyle\pm\mathbf{0.0219}$ $\mathbf{0.3714}\scriptstyle\pm\mathbf{0.0206}$ $0.8444\scriptstyle\pm 0.0385$ $0.8667\scriptstyle\pm 0.0333$
Table 16: Model performances on SKCM

### v2

@@ E.2 Human evaludation and LLM as judge
L3 + Omics $0233\scriptstyle\pm 0.0252 0.0$$020\scriptstyle\pm 0.0020 0.0$$035\scriptstyle\pm 0.0033 0.0$$018\scriptstyle\pm 0.0017 0.0$$133\scriptstyle\pm 0.0231 0.0$$0067\scriptstyle\pm 0.0115$
L3 + Omics + KG $0055\scriptstyle\pm 0.0055 0.0$$033\scriptstyle\pm 0.0031 0.0$$041\scriptstyle\pm 0.0039 0.0$$21\scriptstyle\pm 0.0020 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0300\scriptstyle\pm 0.0249 0.0$$220\scriptstyle\pm 0.0278 0.0$$199\scriptstyle\pm 0.0214 0.0$$104\scriptstyle\pm 0.0114 0.0$$267\scriptstyle\pm 0.0462 0.0$$0333\scriptstyle\pm 0.0306$
L3-FT(Med) + Omics + KG $0044\scriptstyle\pm 0.0038 0.0$$033\scriptstyle\pm 0.0031 0.0$$032\scriptstyle\pm 0.0028 0.0$$16\scriptstyle\pm 0.0014 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5011\scriptstyle\pm 0.0335$ $0.4727\scriptstyle\pm 0.0514$ $0.4862\scriptstyle\pm 0.0430$ $0.3230\scriptstyle\pm 0.0384$ $0.8400\scriptstyle\pm 0.0800$ $0.8267\scriptstyle\pm 0.0808$
L3-FT(CRISPR) + Omics + KG $0.5116\scriptstyle\pm 0.0216$ $0.4807\scriptstyle\pm 0.0469$ $0.4950\scriptstyle\pm 0.0351$ $0.3301\scriptstyle\pm 0.0312$ $0.8800\scriptstyle\pm 0.0000$ $0.8333\scriptstyle\pm 0.0231$
G-Retriever + pre-GAT $0.4417\scriptstyle\pm 0.0005$ $0.3680\scriptstyle\pm 0.0104$ $0.3995\scriptstyle\pm 0.0073$ $0.2504\scriptstyle\pm 0.0062$ $0.7200\scriptstyle\pm 0.0000$ $0.7667\scriptstyle\pm 0.0115$
RoG $0.5184\scriptstyle\pm 0.0462$ $0.4627\scriptstyle\pm 0.0803$ $0.4850\scriptstyle\pm 0.0658$ $0.3254\scriptstyle\pm 0.0519$ $0.8533\scriptstyle\pm 0.0611$ $0.8600\scriptstyle\pm 0.0400$
SubgraphRAG $0.5306\scriptstyle\pm 0.0066$ $0.4627\scriptstyle\pm 0.0854$ $0.4871\scriptstyle\pm 0.0556$ $0.3243\scriptstyle\pm 0.0473$ $\mathbf{0.9333}\scriptstyle\pm\mathbf{0.0231}$ $0.8533\scriptstyle\pm 0.0115$
GNN-RAG $\mathbf{0.5383}\scriptstyle\pm\mathbf{0.0306}$ $0.4627\scriptstyle\pm 0.0463$ $0.4909\scriptstyle\pm 0.0404$ $0.3295\scriptstyle\pm 0.0352$ $0.8533\scriptstyle\pm 0.0462$ $\mathbf{0.8867}\scriptstyle\pm\mathbf{0.0115}$
GALAX $0.5272\scriptstyle\pm 0.0143$ $0.5120\scriptstyle\pm 0.0087$ $\mathbf{0.5192}\scriptstyle\pm\mathbf{0.0103}$ $\mathbf{0.3512}\scriptstyle\pm\mathbf{0.0098}$ $0.9200\scriptstyle\pm 0.0693$ $0.8400\scriptstyle\pm 0.0200$
GALAX (Qwen2.5-7B) $0.5177\scriptstyle\pm 0.0124$ $\mathbf{0.5187}\scriptstyle\pm\mathbf{0.0061}$ $0.5177\scriptstyle\pm 0.0083$ $0.3498\scriptstyle\pm 0.0076$ $0.9200\scriptstyle\pm 0.0800$ $0.8733\scriptstyle\pm 0.0503$
Table 20: Model performance on PAAD
@@
L3 + Omics $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3 + Omics + KG $139\scriptstyle\pm 0.0241 0.000$$08\scriptstyle\pm 0.0014 0.0$$16\scriptstyle\pm 0.0027 0.000$$08\scriptstyle\pm 0.0014 0.0$$167\scriptstyle\pm 0.0289 0.0$$0083\scriptstyle\pm 0.0144$
L3-FT(Med) + Omics $0051\scriptstyle\pm 0.0061 0.0$$050\scriptstyle\pm 0.0043 0.0$$039\scriptstyle\pm 0.0036 0.0$$020\scriptstyle\pm 0.0018 0.0$$167\scriptstyle\pm 0.0289 0.0$$0083\scriptstyle\pm 0.0144$
L3-FT(Med) + Omics + KG $005\scriptstyle\pm 0.0008 0.000$$8\scriptstyle\pm 0.0014 0.000$$6\scriptstyle\pm 0.0010 0.000$$3\scriptstyle\pm 0.0005 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5649\scriptstyle\pm 0.0559$ $0.5358\scriptstyle\pm 0.0686$ $0.5492\scriptstyle\pm 0.0605$ $0.3811\scriptstyle\pm 0.0584$ $0.8000\scriptstyle\pm 0.0500$ $0.8417\scriptstyle\pm 0.0577$
L3-FT(CRISPR) + Omics + KG $0.5499\scriptstyle\pm 0.0570$ $0.5183\scriptstyle\pm 0.0794$ $0.5330\scriptstyle\pm 0.0683$ $0.3657\scriptstyle\pm 0.0650$ $0.8333\scriptstyle\pm 0.0764$ $0.8333\scriptstyle\pm 0.0520$
G-Retriever + pre-GAT $0.5594\scriptstyle\pm 0.0114$ $0.4567\scriptstyle\pm 0.0058$ $0.5016\scriptstyle\pm 0.0009$ $0.3387\scriptstyle\pm 0.0007$ $0.8500\scriptstyle\pm 0.0000$ $0.8500\scriptstyle\pm 0.0000$
RoG $0.5196\scriptstyle\pm 0.1322$ $0.4842\scriptstyle\pm 0.1461$ $0.4973\scriptstyle\pm 0.1406$ $0.3462\scriptstyle\pm 0.1107$ $0.7667\scriptstyle\pm 0.1528$ $0.7833\scriptstyle\pm 0.1181$
[…]
GNN-RAG $0.4923\scriptstyle\pm 0.0762$ $0.4283\scriptstyle\pm 0.0204$ $0.4522\scriptstyle\pm 0.0411$ $0.3068\scriptstyle\pm 0.0248$ $0.7667\scriptstyle\pm 0.1443$ $0.7917\scriptstyle\pm 0.1233$
GALAX $0.5771\scriptstyle\pm 0.0174$ $\mathbf{0.5708}\scriptstyle\pm\mathbf{0.0118}$ $\mathbf{0.5739}\scriptstyle\pm\mathbf{0.0145}$ $\mathbf{0.4044}\scriptstyle\pm\mathbf{0.0154}$ $0.8333\scriptstyle\pm 0.0289$ $0.8500\scriptstyle\pm 0.0000$
GALAX (Qwen2.5-7B) $0.5394\scriptstyle\pm 0.0417$ $0.5383\scriptstyle\pm 0.0427$ $0.5387\scriptstyle\pm 0.0422$ $0.3718\scriptstyle\pm 0.0402$ $0.8167\scriptstyle\pm 0.0289$ $0.8333\scriptstyle\pm 0.0144$
Table 21: Model performance on GBM
@@
L3 + Omics $0056\scriptstyle\pm 0.0096 0.0$$011\scriptstyle\pm 0.0019 0.0$$19\scriptstyle\pm 0.0032 0.000$$9\scriptstyle\pm 0.0016 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + Omics + KG $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$ $\scriptstyle\pm 0.0000 0.0000±0.0000$$\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0069\scriptstyle\pm 0.0045 0.0$$133\scriptstyle\pm 0.0088 0.0$$090\scriptstyle\pm 0.0059 0.0$$46\scriptstyle\pm 0.0030 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0107\scriptstyle\pm 0.0158 0.0$$056\scriptstyle\pm 0.0038 0.0$$064\scriptstyle\pm 0.0074 0.0$$32\scriptstyle\pm 0.0037 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(CRISPR) + Omics $0.5641\scriptstyle\pm 0.0310$ $0.5333\scriptstyle\pm 0.0462$ $0.5480\scriptstyle\pm 0.0387$ $0.3794\scriptstyle\pm 0.0364$ $0.9333\scriptstyle\pm 0.0667$ $0.9000\scriptstyle\pm 0.0333$
L3-FT(CRISPR) + Omics + KG $0.5391\scriptstyle\pm 0.0213$ $0.5178\scriptstyle\pm 0.0164$ $0.5281\scriptstyle\pm 0.0181$ $0.3596\scriptstyle\pm 0.0165$ $\mathbf{0.9556}\scriptstyle\pm\mathbf{0.0385}$ $0.8444\scriptstyle\pm 0.0694$
G-Retriever + pre-GAT $0.4906\scriptstyle\pm 0.0075$ $0.4411\scriptstyle\pm 0.0038$ $0.4642\scriptstyle\pm 0.0010$ $0.3037\scriptstyle\pm 0.0000$ $\mathbf{0.9556}\scriptstyle\pm\mathbf{0.0385}$ $0.8333\scriptstyle\pm 0.0577$
RoG $\mathbf{0.5910}\scriptstyle\pm\mathbf{0.0281}$ $0.5078\scriptstyle\pm 0.1165$ $0.5380\scriptstyle\pm 0.0853$ $0.3734\scriptstyle\pm 0.0756$ $\mathbf{0.9556}\scriptstyle\pm\mathbf{0.0770}$ $\mathbf{0.9444}\scriptstyle\pm\mathbf{0.0509}$
SubgraphRAG $0.5066\scriptstyle\pm 0.0597$ $0.4544\scriptstyle\pm 0.1001$ $0.4707\scriptstyle\pm 0.0884$ $0.3176\scriptstyle\pm 0.0729$ $0.7778\scriptstyle\pm 0.1925$ $0.7556\scriptstyle\pm 0.2117$
GNN-RAG $0.5567\scriptstyle\pm 0.0222$ $0.5478\scriptstyle\pm 0.0212$ $0.5521\scriptstyle\pm 0.0219$ $0.3855\scriptstyle\pm 0.0141$ $0.8667\scriptstyle\pm 0.2309$ $0.8111\scriptstyle\pm 0.1540$
GALAX $0.5527\scriptstyle\pm 0.0021$ $0.5422\scriptstyle\pm 0.0139$ $0.5473\scriptstyle\pm 0.0081$ $0.3769\scriptstyle\pm 0.0079$ $0.9111\scriptstyle\pm 0.0770$ $0.8667\scriptstyle\pm 0.0000$
GALAX (Qwen2.5-7B) $0.5715\scriptstyle\pm 0.0228$ $\mathbf{0.5656}\scriptstyle\pm\mathbf{0.0269}$ $\mathbf{0.5685}\scriptstyle\pm\mathbf{0.0249}$ $\mathbf{0.3985}\scriptstyle\pm\mathbf{0.0236}$ $0.9111\scriptstyle\pm 0.0770$ $0.8556\scriptstyle\pm 0.0385$
Table 22: Model performance on SARC
@@
L3 + Omics $0111\scriptstyle\pm 0.0192 0.0$$033\scriptstyle\pm 0.0058 0.0$$051\scriptstyle\pm 0.0089 0.0$$26\scriptstyle\pm 0.0045 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
@@
L3-FT(Med) + Omics $0689\scriptstyle\pm 0.0631 0.0$$167\scriptstyle\pm 0.0115 0.0$$229\scriptstyle\pm 0.0180 0.0$$119\scriptstyle\pm 0.0094 0.0$$222\scriptstyle\pm 0.0385 0.0$$0667\scriptstyle\pm 0.0577$
@@
L3-FT(CRISPR) + Omics $0.5347\scriptstyle\pm 0.0167$ $0.5089\scriptstyle\pm 0.0353$ $0.5211\scriptstyle\pm 0.0266$ $0.3547\scriptstyle\pm 0.0253$ $\mathbf{0.9111}\scriptstyle\pm\mathbf{0.0385}$ $0.8778\scriptstyle\pm 0.0192$
L3-FT(CRISPR) + Omics + KG $0.5423\scriptstyle\pm 0.0207$ $0.5022\scriptstyle\pm 0.0539$ $0.5207\scriptstyle\pm 0.0383$ $0.3570\scriptstyle\pm 0.0344$ $0.8444\scriptstyle\pm 0.0385$ $0.8444\scriptstyle\pm 0.0385$
G-Retriever + pre-GAT $0.4461\scriptstyle\pm 0.0308$ $0.3667\scriptstyle\pm 0.0115$ $0.4009\scriptstyle\pm 0.0055$ $0.2518\scriptstyle\pm 0.0051$ $0.8667\scriptstyle\pm 0.0000$ $\mathbf{0.9222}\scriptstyle\pm\mathbf{0.0192}$
RoG $\mathbf{0.5533}\scriptstyle\pm\mathbf{0.0201}$ $0.4589\scriptstyle\pm 0.0740$ $0.4922\scriptstyle\pm 0.0433$ $0.3324\scriptstyle\pm 0.0375$ $0.8889\scriptstyle\pm 0.0385$ $0.8778\scriptstyle\pm 0.0192$
SubgraphRAG $0.5365\scriptstyle\pm 0.0277$ $0.4456\scriptstyle\pm 0.0168$ $0.4772\scriptstyle\pm 0.0188$ $0.3175\scriptstyle\pm 0.0166$ $0.8889\scriptstyle\pm 0.1018$ $0.8667\scriptstyle\pm 0.0882$
GNN-RAG $0.5485\scriptstyle\pm 0.0097$ $0.4778\scriptstyle\pm 0.0587$ $0.5036\scriptstyle\pm 0.0423$ $0.3426\scriptstyle\pm 0.0403$ $0.8889\scriptstyle\pm 0.0385$ $0.8889\scriptstyle\pm 0.0385$
GALAX $0.5414\scriptstyle\pm 0.0143$ $0.5144\scriptstyle\pm 0.0168$ $0.5271\scriptstyle\pm 0.0157$ $0.3602\scriptstyle\pm 0.0165$ $\mathbf{0.9111}\scriptstyle\pm\mathbf{0.1018}$ $0.8444\scriptstyle\pm 0.0509$
GALAX (Qwen2.5-7B) $0.5383\scriptstyle\pm 0.0096$ $\mathbf{0.5311}\scriptstyle\pm\mathbf{0.0084}$ $\mathbf{0.5347}\scriptstyle\pm\mathbf{0.0090}$ $\mathbf{0.3674}\scriptstyle\pm\mathbf{0.0088}$ $\mathbf{0.9111}\scriptstyle\pm\mathbf{0.1018}$ $0.8889\scriptstyle\pm 0.0694$
Table 23: Model performance on OV
@@
L3 + Omics $0241\scriptstyle\pm 0.0251 0.0$$044\scriptstyle\pm 0.0051 0.0$$075\scriptstyle\pm 0.0085 0.0$$38\scriptstyle\pm 0.0043 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3 + Omics + KG $0660\scriptstyle\pm 0.0667 0.0$$111\scriptstyle\pm 0.0102 0.0$$183\scriptstyle\pm 0.0161 0.0$$93\scriptstyle\pm 0.0082 0.000$$00\scriptstyle\pm 0.0000 0.0$$0889\scriptstyle\pm 0.0770$
L3-FT(Med) + Omics $0356\scriptstyle\pm 0.0412 0.0$$100\scriptstyle\pm 0.0067 0.0$$102\scriptstyle\pm 0.0072 0.0$$051\scriptstyle\pm 0.0036 0.0$$444\scriptstyle\pm 0.0770 0.0$$0222\scriptstyle\pm 0.0385$
L3-FT(Med) + Omics + KG $0797\scriptstyle\pm 0.0496 0.0$$344\scriptstyle\pm 0.0280 0.0$$457\scriptstyle\pm 0.0375 0.0$$253\scriptstyle\pm 0.0226 0.0$$000\scriptstyle\pm 0.0000 0.0$$0333\scriptstyle\pm 0.0333$
L3-FT(CRISPR) + Omics $0.5469\scriptstyle\pm 0.0274$ $0.5167\scriptstyle\pm 0.0406$ $0.5307\scriptstyle\pm 0.0325$ $0.3631\scriptstyle\pm 0.0295$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $0.9111\scriptstyle\pm 0.0694$
L3-FT(CRISPR) + Omics + KG $0.5439\scriptstyle\pm 0.0066$ $0.4989\scriptstyle\pm 0.0195$ $0.5200\scriptstyle\pm 0.0123$ $0.3542\scriptstyle\pm 0.0095$ $0.9556\scriptstyle\pm 0.0385$ $0.9444\scriptstyle\pm 0.0385$
G-Retriever + pre-GAT $0.5413\scriptstyle\pm 0.0525$ $0.4033\scriptstyle\pm 0.0173$ $0.4575\scriptstyle\pm 0.0067$ $0.2991\scriptstyle\pm 0.0041$ $\mathbf{1.0000}\scriptstyle\pm\mathbf{0.0000}$ $\mathbf{0.9667}\scriptstyle\pm\mathbf{0.0000}$
RoG $0.5708\scriptstyle\pm 0.0158$ $0.5289\scriptstyle\pm 0.0523$ $0.5453\scriptstyle\pm 0.0343$ $0.3774\scriptstyle\pm 0.0317$ $0.9556\scriptstyle\pm 0.0385$ $0.9222\scriptstyle\pm 0.0385$
SubgraphRAG $\mathbf{0.6143}\scriptstyle\pm\mathbf{0.0414}$ $0.4833\scriptstyle\pm 0.0384$ $0.5289\scriptstyle\pm 0.0215$ $0.3650\scriptstyle\pm 0.0185$ $0.9333\scriptstyle\pm 0.0667$ $0.9444\scriptstyle\pm 0.0385$
GNN-RAG $0.5687\scriptstyle\pm 0.0300$ $0.4600\scriptstyle\pm 0.0145$ $0.5026\scriptstyle\pm 0.0061$ $0.3380\scriptstyle\pm 0.0036$ $0.8444\scriptstyle\pm 0.1540$ $0.8111\scriptstyle\pm 0.1540$
GALAX $0.5489\scriptstyle\pm 0.0176$ $0.5344\scriptstyle\pm 0.0259$ $0.5413\scriptstyle\pm 0.0219$ $0.3714\scriptstyle\pm 0.0206$ $0.8444\scriptstyle\pm 0.0385$ $0.8667\scriptstyle\pm 0.0333$
GALAX (Qwen2.5-7B) $0.5643\scriptstyle\pm 0.0190$ $\mathbf{0.5689}\scriptstyle\pm\mathbf{0.0550}$ $\mathbf{0.5654}\scriptstyle\pm\mathbf{0.0355}$ $\mathbf{0.3957}\scriptstyle\pm\mathbf{0.0349}$ $0.9778\scriptstyle\pm 0.0385$ $0.9556\scriptstyle\pm 0.0192$
Table 24: Model performance on SKCM